# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws

### v1

@@ V-B Feasibility of the Disturbed Bang-Singular-Bang Trajectory
The feasibility of the disturbed trajectory is proved by two steps. Firstly, the consistent error between the disturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory. In this way, the disturbed trajectory can be proved feasible at the points far from keypoints. Secondly, the feasibility of the disturbed trajectory near the keypoints is proved based on Proposition 3, Proposition 4, and Proposition 5. As a result, the feasibility of the disturbed trajectory is proved. The two steps correspond to Proposition 6 and Theorem 2, respectively.
@@
${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is driven by $\forall i\in\left[M\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$. $\exists C>0$, s.t. $\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$,
$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon,$ (30a)
$\max_{i\in\left[M\right]}\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|\leq C\varepsilon.$ (30b)
Among them, $\forall i\in\left[M\right]$, $t\in\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$. ${\boldsymbol{x}}^{\prime}\left(t_{0}\right)={\boldsymbol{x}}\left(t_{0}\right)$ and $t_{0}^{\prime}=t_{0}$. ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$ and ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$. $0<\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$.
@@
Since $\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$ and $\left|t_{i}-t_{i}^{\prime}\right|\leq\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, it has $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$. $\forall i\in\left[M\right]$, let $\hat{t}_{2i-1}=\min\left\{t_{i},t_{i}^{\prime}\right\}$, $\hat{t}_{2i}=\max\left\{t_{i},t_{i}^{\prime}\right\}$, and $\hat{t}_{0}=0$. Then $\left\{\hat{t}_{i}\right\}_{i=0}^{2M}$ increases monotonically.
@@
Therefore,
$\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (32)
$=$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\left({\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right)\right\|$
$\leq$ $\sup_{t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$
@@
Therefore,
$\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (35)
$\leq$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\varepsilon\sup_{t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]}\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j^{\prime}}\right\|$
@@
Therefore, (30a) holds. Similarly, $\forall i\in\left[M\right]$,
$\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|$ (37)
$\leq$ $\left\|{\boldsymbol{x}}\left(t_{i}^{\prime}\right)-{\boldsymbol{x}}^{\prime}\left(t_{i}^{\prime}\right)\right\|+\left\|{\boldsymbol{x}}\left(t_{i}\right)-{\boldsymbol{x}}\left(t_{i}^{\prime}\right)\right\|$
$\leq$ $\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}+C_{1}\left\|{\boldsymbol{x}}\right\|\varepsilon+C_{2}C_{3}\varepsilon$
$\leq$ $\left(2C_{3}^{2M-2}-1\right)\left(C_{1}\left\|{\boldsymbol{x}}\right\|+C_{2}C_{3}\right)\varepsilon.$
Therefore, (30b) holds. ∎
@@
${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are Bang-Singular-Bang trajectories of problem (1a) without chattering. Assume that ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ are represented by the same augmented switching law ${\mathcal{L}}={\mathcal{L}}^{\prime}$. ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ is feasible for problem (1a). $\exists\varepsilon>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a).
@@
According to Proposition 6, $\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then (30a) holds.
@@
Secondly, consider the inequality constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\leq 0$. Denote that $\left\{i^{\left(p\right)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t. $i^{\left(p\right)}_{0}=0$, $i^{\left(p\right)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds. If ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}},t_{i^{\left(p\right)}_{j}}\right)$, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{\left(p\right)}_{j-1}}^{\prime},t_{i^{\left(p\right)}_{j}}^{\prime}\right)$ since ${\mathcal{L}}={\mathcal{L}}^{\prime}$.

### v2

@@ V-B Feasibility of the Perturbed BBS Trajectory
The feasibility of the perturbed trajectory is proved by two steps. Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory. In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory. Secondly, Theorem 2 proves the perturbed trajectory’s feasibility based on Propositions 4, 5, 6, and 7. In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible. Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
@@
Consider a BBS trajectory ${\boldsymbol{x}}\left(t\right)$ satisfying $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$. Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$. $\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. $\forall\varepsilon\in\left(0,\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}\right)$, $\forall{\boldsymbol{x}}^{\prime}\left(t\right)$ is a perturbed trajectory induced by ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ with $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$,
$\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}\triangleq\sup_{t\in\left[t_{0},\min\left\{t_{M},t_{M}^{\prime}\right\}\right]}\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq C\varepsilon$ (30)
where $\forall i\in\left[M\right]$, $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{i}{\boldsymbol{x}}^{\prime}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1}^{\prime},t_{i}^{\prime}\right)$ and $t_{0}^{\prime}=t_{0}$.
@@
Since $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon<\frac{1}{2}\min_{i\in\left[M-1\right]}\left\{\left|t_{i+1}-t_{i}\right|\right\}$, it holds that $\forall i\in\left[M-1\right]$, $\max\left\{t_{i},t_{i}^{\prime}\right\}<\min\left\{t_{i+1},t_{i+1}^{\prime}\right\}$. $\forall i\in\left[M\right]$, let $\hat{t}_{2i-1}=\min\left\{t_{i},t_{i}^{\prime}\right\}$, $\hat{t}_{2i}=\max\left\{t_{i},t_{i}^{\prime}\right\}$, and $\hat{t}_{0}=t_{0}$. Then $\left\{\hat{t}_{i}\right\}_{i=0}^{2M}$ increases monotonically.
@@
Therefore, $\forall t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$, it holds that
$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$ (32)
@@
Therefore, $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$, it holds that
$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|$ (35)
$\leq$ $\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i-1}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i-1}\right)\right\|$
$+$ $\varepsilon\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{j}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j}-{\mathrm{e}}^{{\boldsymbol{A}}_{j^{\prime}}\left(t-\hat{t}_{2i-1}\right)}{\boldsymbol{b}}_{j^{\prime}}\right\|$
@@
Therefore, (30) holds.
@@
Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is ${\mathcal{L}}$ in problem (1a). Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. $\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is trajectory also represented by ${\mathcal{L}}^{\prime}={\mathcal{L}}$, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a). ${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by ${\boldsymbol{t}}$ and ${\boldsymbol{t}}^{\prime}$ in the same way as that in Proposition 7.
@@
According to Proposition 7, $\exists C>0$ only dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then (30) holds.
@@
Secondly, consider the inequality constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\leq 0$. Denote $\left\{i^{(p)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t. $i^{(p)}_{0}=0$, $i^{(p)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ switches between active and inactive at $t_{i^{(p)}_{j}}$. In other words, $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i_{j}^{(p)}}\right)+d_{p}=0$ holds. Furthermore, $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds. Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, the above property on $\left({\boldsymbol{x}}\left(t\right),{\boldsymbol{t}}\right)$ holds for $\left({\boldsymbol{x}}^{\prime}\left(t\right),{\boldsymbol{t}}^{\prime}\right)$ as well.
Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig. 3(a). ${\mathcal{L}}={\mathcal{L}}^{\prime}$ implies that ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\equiv 0$ holds. Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right)$.
Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}<0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig. 3(b). The feasibility of ${\boldsymbol{x}}^{\prime}\left(t\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right]$ is discussed in two parts, i.e., far from and near the keypoints.